# Fair Resource Allocation in Weakly Coupled Markov Decision Processes

### v1

@@ Abstract
We consider fair resource allocation in sequential decision-making environments modeled as weakly coupled Markov decision processes, where resource constraints couple the action spaces of $N$ sub-Markov decision processes (sub-MDPs) that would otherwise operate independently. We adopt a fairness definition using the generalized Gini function instead of the traditional utilitarian (total-sum) objective. After introducing a general but computationally prohibitive solution scheme based on linear programming, we focus on the homogeneous case where all sub-MDPs are identical. For this case, we show for the first time that the problem reduces to optimizing the utilitarian objective over the class of “permutation invariant” policies. This result is particularly useful as we can exploit Whittle index policies in the restless bandits setting while, for the more general setting, we introduce a count-proportion-based deep reinforcement learning approach. Finally, we validate our theoretical findings with comprehensive experiments, confirming the effectiveness of our proposed method in achieving fairness.
@@ 1 Introduction
Machine learning (ML) algorithms play a significant role in automated decision-making processes, influencing our daily lives. Mitigating biases within the ML pipeline is crucial to ensure fairness and generate reliable outcomes (Caton and Haas,, 2024). Extensive research has been conducted to enhance fairness across various applications, such as providing job matching services (van den Broek et al.,, 2020; Raghavan et al.,, 2020), assigning credit scores and loans (Kozodoi et al.,, 2022), and delivering healthcare services (Farnadi et al.,, 2021; Chen et al.,, 2023).
However, most real-world decision processes are sequential in nature and past decisions may have an impact on equity. For example, if people are unfairly denied credit or job opportunities early in their careers, it can have long-term consequences for their financial well-being and opportunities for advancement (Liu et al.,, 2018). In addition, the impact of decisions can accumulate, which may lead to large disparities over time and can actually exacerbate unfairness (Creager et al.,, 2020; D’Amour et al.,, 2020).
Another motivating example for this work is taxi dispatching: when certain areas are consistently prioritized over others, it can lead to long-term disparities in service accessibility. This may lead to long waiting times for passengers in certain neighborhoods, while taxis run empty and seek passengers in other areas (Liu et al.,, 2021). Thus, to ensure that individuals are not treated unfairly based on past decisions or changing circumstances, we should extend the current definitions of fairness to consider the impact of feedback effects (Ghalme et al., 2022a, ).
Fairness is a complex and multi-faceted concept, and there are many different ways in which it can be operationalized and measured. We resort to the generalized Gini social welfare function (GGF) (Weymark,, 1981), which covers various fairness measures as special cases. The long-term impacts of fair decision dynamics have recently been approached using Markov decision processes (MDPs) (Wen et al.,, 2021; Puranik et al.,, 2022; Ghalme et al., 2022b, ). By studying fairness in MDPs, we can better understand how such considerations can be incorporated into decision-making processes, how they may affect the outcomes of those processes, and how to develop approaches to promote fairness in a wider range of settings.
To the best of our knowledge, we are the first to incorporate fairness considerations in the form of the GGF objective within weakly coupled Markov decision processes (WCMDPs) (Hawkins,, 2003; Adelman and Mersereau,, 2008), which can also be considered as multi-action multi-resource restless multi-arm bandits. This model captures the complex interactions among coupled MDPs over time and allows the applicability of our work to various applications in scheduling (Saure et al.,, 2012; El Shar and Jiang,, 2024), application screening (Gast et al.,, 2024), budget allocation (Boutilier and Lu,, 2016), inventory (El Shar and Jiang,, 2024), and restless multi-arm bandits (RMABs) (Hawkins,, 2003; Zhang,, 2022).
@@
There are two studies closely related to our work. The first work by Gast et al., (2024) considers symmetry simplification and count aggregation MDPs. They focus on solving a LP model repeatedly with a total-sum objective to obtain asymptotic optimal solutions when the number of coupled MDPs is very large, whereas we explicitly address the fairness aspect and exploit a state count representation to design scalable deep RL approaches. The second related work by Siddique et al., (2020) integrates the fair Gini multi-objective RL to treat every user equitably. This fair optimization problem is later extended to the decentralized cooperative multi-agent RL by Zimmer et al., (2021), and further refined to incorporate preferential treatment with human feedback by Siddique et al., (2023) and Yu et al., (2023). In contrast, our work demonstrates that the WCMDP with the GGF objective and identical coupled MDPs reduces to a much simpler utilitarian problem, which allows us to exploit its structure to develop efficient and scalable algorithms. A more comprehensive literature review on fairness in MDPs, RL, and RMABs, is provided in Appendix A to clearly position our work.
@@ 2 Background
We review infinite-horizon WCMDPs and introduce the GGF for encoding fairness. We then define the fair optimization problem and provide an exact solution scheme based on linear programming.
@@ 2.1 The Weakly Coupled MDP
An infinite-horizon WCMDP $\mathcal{M}^{(N)}$ consists of $N$ sub-MDPs, where each sub-MDP is independent of the others in terms of state transitions and rewards. They are linked to each other solely through a set of $K$ constraints on their actions at each time step. Formally, the WCMDP is defined by a tuple $(\mathcal{S}^{(N)},\mathcal{A}^{(N)},p^{(N)},\bm{r},\mu^{(N)},\gamma)$, where the state space $\mathcal{S}^{(N)}$ is the Cartesian product of individual state spaces, and the action space $\mathcal{A}^{(N)}$ is a subset of the Cartesian product of action spaces, defined as $\mathcal{A}^{(N)}:=\{(a_{1},\dots,a_{N})\mid\sum_{n=1}^{N}d_{k,n}(a_{n})\leq b_{k},\forall k\in{\mathcal{K}},\ a_{n}\in\mathcal{A}_{n},\forall n\in{\mathcal{N}}\},$ where ${\mathcal{K}}:=[K]$ is the index set of constraints, $d_{k,n}(a_{n})\in{\mathbb{R}}_{+}$ represents the consumption of the $k$-th resource consumption by the $n$-th MDP when action $a_{n}$ is taken, and $b_{k}\in\mathbb{R}_{+}$ the available resource of type $k$ [^1] Actually, $b_{k}\leq\sum_{n=1}^{N}\max_{a_{n}\in\mathcal{A}_{n}}d_{k,n}(a_{n})$, w.l.o.g. We define an idle action that consumes no resources for any resource $k$ to ensure that the feasible action space is non-empty.
The state transitions of the sub-MDPs are independent, so the system transits from state $\bm{s}$ to state $\bm{s}^{\prime}$ for a given feasible action $\bm{a}$ at time $t$ with probability $p^{(N)}(\bm{s}^{\prime}|\bm{s},\bm{a})=\prod^{N}_{n=1}p_{n}(s^{\prime}_{n}|s_{n},a_{n})=\prod^{N}_{n=1}\mathbb{P}(s_{t+1,n}=s^{\prime}_{n}|s_{t,n}=s_{n},a_{t,n}=a_{n})$. After choosing an action $\bm{a}\in\mathcal{A}^{(N)}$ in the state $\bm{s}\in\mathcal{S}^{(N)}$, the decision maker receives rewards defined as $\bm{r}(\bm{s},\bm{a})=(r_{1}(s_{1},a_{1}),\dots,r_{N}(s_{N},a_{N}))$ with each component representing the reward associated with the respective sub-MDP $\mathcal{M}_{n}$. We employ a vector form for the rewards to offer the flexibility for formulating fairness objectives on individual expected total discounted rewards in later sections.
@@ 2.2 The Generalized Gini Function
The vector $\bm{V}^{\pi}_{0}$ represents expected utilities for users. A social welfare function aggregates these utilities into a scalar, measuring fairness in utility distribution with respect to a maximization objective.
Social welfare functions can vary depending on the values of a society, such as $\alpha$-fairness (Mo and Walrand,, 2000) or max-min fairness (Bistritz et al.,, 2020). Following Siddique et al., (2020), we require a fair solution to meet three properties: efficiency, impartiality, and equity. This motivates the use of the GGF from economics (Weymark,, 1981), which satisfies these properties. For $N$ users, the GGF is defined as $\operatorname{GGF}_{{\bm{w}}}[{\bm{v}}]:=\min_{\sigma\in{\mathbb{S}}^{N}}\sum_{n=1}^{N}{w}_{n}{v}_{\sigma(n)}$, where ${\bm{v}}\in{\mathbb{R}}^{N}$, ${\bm{w}}\in\Delta({\mathcal{N}})$ is non-increasing in $n$, i.e., $w_{1}\geq w_{2}\geq\cdots\geq w_{N}$. Intuitively, since $\operatorname{GGF}_{{\bm{w}}}[{\bm{v}}]=\sum_{n=1}^{N}{w}_{n}{v}_{\sigma^{*}(n)}$ with $\sigma^{*}$ as the minimizer, which reorders the terms of ${\bm{v}}$ from lowest to largest, it computes the weighted sum of ${\bm{v}}$ assigning larger weights to its lowest components.
As discussed in Siddique et al., (2020), the GGF can reduce to special cases by setting the weights to specific values, including the maxmin egalitarian approach ($w_{1}\rightarrow 1,w_{2}\rightarrow 0,\ldots,w_{N}\rightarrow 0$) (Rawls,, 1971), regularized maxmin egalitarian ($w_{1}\rightarrow 1,w_{2}\rightarrow\epsilon,\ldots,w_{N}\rightarrow\epsilon$), leximin notion of fairness ($w_{k}/w_{k+1}\rightarrow\infty$) (Rawls,, 1971; Moulin,, 1991), and the utilitarian approach formally defined below for the later use in reducing the GGF problem.
@@
The utilitarian approach maximizes the average utilities over all individuals but does not guarantee fairness in utility distribution, as some users may be disadvantaged to increase overall utility. The GGF offers flexibility by encoding various fairness criteria in a structured way. Moreover, $\operatorname{GGF}_{{\bm{w}}}[{\bm{v}}]$ is concave in ${\bm{v}}$, which has nice properties for problem reformulation.
@@ 2.3 The GGF-WCMDP Problem
We note that Lemma 3.1 in (Siddique et al.,, 2020) establishes the optimality of stationary Markov policies for any multi-objective discounted infinite-horizon MDP under the GGF criterion. The optimal policy for the GGF-WCMDP problem (2) can be computed by solving the following LP model with the GGF objective (GGF-LP):
@@
$\mathrm{s.t.}$ $\lambda_{i}+\nu_{j}\leq w_{i}\sum\limits_{\bm{s}\in\mathcal{S}^{(N)}}\sum\limits_{\bm{a}\in\mathcal{A}^{(N)}}\bm{r}_{j}(\bm{s},\bm{a})q(\bm{s},\bm{a}),\quad\forall i,j{\in{\mathcal{N}}},$ (3d)
$\sum\limits_{\bm{a}\in\mathcal{A}^{(N)}}q(\bm{s},\bm{a})-\gamma\sum\limits_{\bm{s}^{\prime}\in\mathcal{S}^{(N)}}\sum\limits_{\bm{a}\in\mathcal{A}^{(N)}}q(\bm{s}^{\prime},\bm{a})p^{(N)}(\bm{s}|\bm{s}^{\prime},\bm{a})=\mu^{(N)}(\bm{s}),\quad\forall\bm{s}\in\mathcal{S}^{(N)},$
$q(\bm{s},\bm{a})\geq 0,\quad\forall\bm{s}\in\mathcal{S}^{(N)},\quad\forall\bm{a}\in\mathcal{A}^{(N)}.$
@@
The dual form separates dynamics from rewards, with the expected discounted reward for sub-MDP $n$ given by $\sum_{\bm{s}\in\mathcal{S}^{(N)}}\sum_{\bm{a}\in\mathcal{A}^{(N)}}\bm{r}_{n}(\bm{s},\bm{a})q(\bm{s},\bm{a})$. The one-to-one mapping between the solution $q(\bm{s},\bm{a})$ and an optimal policy $\pi(\bm{s},\bm{a})$ is $\pi(\bm{s},\bm{a})=q(\bm{s},\bm{a})/\sum_{\bm{a}\in\mathcal{A}^{(N)}}q(\bm{s},\bm{a})$.
Scalability is a critical challenge in obtaining exact solutions as the state and action spaces grow exponentially with respect to the number of sub-MDPs, making the problem intractable. We thus explore approaches that exploit symmetric problem structures, apply count-based state aggregation, and use RL-based approximation methods, to address this scalability issue, which will be discussed next.
@@ A WCMDP is symmetric if
This symmetry ensures that the expected state-value function, when averaged over all trajectories, is identical for each sub-MDP, leading to a uniform state-value representation. From this observation, and applying Theorem 6.9.1 from (Puterman,, 2005), we construct a permutation-invariant policy from any policy, resulting in uniform state-value (Lemma 3.3).
@@
This theorem simplifies solving the GGF-WCMDP problem by reducing it to an equivalent utilitarian problem, showing that at least one permutation-invariant policy is optimal for the original GGF-WCMDP problem and the utilitarian reduction. Therefore, we can restrict the search for optimal policies to this specific class of permutation-invariant policies. The utilitarian approach does not compromise the GGF optimality and allows us to leverage more efficient and scalable techniques to solve the GGF-WCMDP problem (Eq. 2), such as the Whittle index policies for RMABs, as demonstrated in the experimental section.

### v2

@@ Abstract
We consider fair resource allocation in sequential decision-making environments modeled as weakly coupled Markov decision processes, where resource constraints couple the action spaces of $N$ sub-Markov decision processes (sub-MDPs) that would otherwise operate independently. We adopt a fairness definition using the generalized Gini function instead of the traditional utilitarian (total-sum) objective. After introducing a general but computationally prohibitive solution scheme based on linear programming, we focus on the homogeneous case where all sub-MDPs are identical. For this case, we show for the first time that the problem reduces to optimizing the utilitarian objective over the class of “permutation invariant” policies. This result is particularly useful as we can exploit efficient algorithms that optimizes the utilitarian objective such as Whittle index policies in restless bandits to solve the problem with this fairness objective. For more general settings, we introduce a count-proportion-based deep reinforcement learning approach. Finally, we validate our theoretical findings with comprehensive experiments, confirming the effectiveness of our proposed method in achieving fairness.
@@ 1 INTRODUCTION
Machine learning (ML) algorithms play a significant role in automated decision-making processes, influencing our daily lives. Mitigating biases within the ML pipeline is crucial to ensure fairness and generate reliable outcomes (Caton and Haas,, 2024). Extensive research has been conducted to enhance fairness across various applications, such as providing job hiring services (van den Broek et al.,, 2020; Cimpean et al.,, 2024), assigning credit scores and loans (Kozodoi et al.,, 2022), and delivering healthcare services (Farnadi et al.,, 2021; Chen et al.,, 2023).
However, most real-world decision processes are sequential in nature and past decisions may have a long-term impact on equity (D’Amour et al.,, 2020). For example, if people are unfairly denied credit or job opportunities early in their careers, there would be long-term consequences on opportunities for advancement (Liu et al.,, 2018). Another motivating example is taxi dispatching. If certain areas are consistently prioritized over others, then there can be long-term disparities in service accessibility. This may lead to long waiting times for passengers in certain neighborhoods, while taxis run empty and seek passengers in other areas (Liu et al.,, 2021; Guo et al.,, 2023).
Fairness is a complex and multi-faceted concept, and there are many different ways in which it can be operationalized and measured. We resort to the generalized Gini social welfare function (GGF) (Weymark,, 1981), which covers various fairness measures as special cases. The long-term impacts of fair decision dynamics have recently been approached using Markov decision processes (MDPs) (Wen et al.,, 2021; Puranik et al.,, 2022; Ghalme et al.,, 2022). Studying fairness in MDPs helps mitigate bias and inequality in decision-making processes and evaluate their broader societal and operational impacts across diverse applications.
To the best of our knowledge, we are the first to incorporate fairness considerations in the form of the GGF objective within weakly coupled Markov decision processes (WCMDPs) (Hawkins,, 2003; Adelman and Mersereau,, 2008), which can be considered as an extension of restless multi-arm bandit problems (RMABs) (Hawkins,, 2003; Zhang,, 2022) to multi-action and multi-resource settings. This model is particularly relevant to resource allocation problems, as it captures the complex interactions of coupled MDPs (arms) over time restricted by limited resource availability, and allows the applicability of our work to various applications in scheduling (Saure et al.,, 2012; El Shar and Jiang,, 2024), application screening (Gast et al.,, 2024), budget allocation (Boutilier and Lu,, 2016), and inventory (El Shar and Jiang,, 2024).
@@
There are two studies closely related to our work. The first work by Gast et al., (2024) considers symmetry simplification and count aggregation MDPs. They focus on solving an LP model repeatedly with a total-sum objective to obtain asymptotic optimal solutions when the number of coupled MDPs is very large, whereas we explicitly address the fairness aspect and exploit a state count representation to design scalable deep RL approaches. The second work by Siddique et al., (2020) integrates the fair Gini multi-objective RL to treat every user equitably. This fair optimization problem is later extended to the decentralized cooperative multi-agent RL by Zimmer et al., (2021), and further refined to incorporate preferential treatment with human feedback by Siddique et al., (2023) and Yu et al., (2023). In contrast, our work demonstrates that the WCMDP with the GGF objective and identical coupled MDPs reduces to a much simpler utilitarian problem, which allows us to exploit its structure to develop efficient and scalable algorithms. A more comprehensive literature review on fairness in resource allocation, MDPs, RL, and RMABs, is provided in Appendix A to clearly position our work.
@@ 2 BACKGROUND
We start by reviewing infinite-horizon WCMDPs and introducing the GGF for encoding fairness. We then define the fair optimization problem and provide an exact solution scheme based on linear programming.
@@ 2.1 The Weakly Coupled MDP
An infinite-horizon WCMDP $\mathcal{M}^{(N)}$ consists of $N$ sub-MDPs, where each sub-MDP is independent of the others in terms of state transitions and rewards. They are linked to each other solely through a set of $K$ constraints on their actions at each time step. Formally, the WCMDP is defined by a tuple $(\mathcal{S}^{(N)},\mathcal{A}^{(N)},p^{(N)},\bm{r},\mu^{(N)},\gamma)$, where the state space $\mathcal{S}^{(N)}$ is the Cartesian product of individual state spaces, and the action space $\mathcal{A}^{(N)}$ is a subset of the Cartesian product of action spaces, defined as $\mathcal{A}^{(N)}:=\{(a_{1},\dots,a_{N})\mid\sum_{n=1}^{N}d_{k,n}(a_{n})\leq b_{k},\forall k\in{\mathcal{K}},\ a_{n}\in\mathcal{A}_{n}\},$ where ${\mathcal{K}}:=[K]$ is the index set of constraints, $d_{k,n}(a_{n})\in{\mathbb{R}}_{+}$ represents the consumption of the $k$-th resource consumption by the $n$-th MDP when action $a_{n}$ is taken, and $b_{k}\in\mathbb{R}_{+}$ the available resource of type $k$ [^1] Actually, $b_{k}\leq\sum_{n=1}^{N}\max_{a_{n}\in\mathcal{A}_{n}}d_{k,n}(a_{n})$, w.l.o.g. We define an idle action that consumes no resources for any resource $k$ to ensure that the feasible action space is non-empty.
The state transitions of the sub-MDPs are independent, so the system transits from state $\bm{s}$ to state $\bm{s}^{\prime}$ for a given feasible action $\bm{a}$ at time $t$ with probability $p^{(N)}(\bm{s}^{\prime}|\bm{s},\bm{a})=\prod^{N}_{n=1}p_{n}(s^{\prime}_{n}|s_{n},a_{n})=\prod^{N}_{n=1}\mathbb{P}(s_{t+1,n}=s^{\prime}_{n}|s_{t,n}=s_{n},a_{t,n}=a_{n})$. After choosing an action $\bm{a}\in\mathcal{A}^{(N)}$ in state $\bm{s}\in\mathcal{S}^{(N)}$, the decision maker receives rewards defined as $\bm{r}(\bm{s},\bm{a})=(r_{1}(s_{1},a_{1}),\dots,r_{N}(s_{N},a_{N}))$ with each component representing the reward associated with the respective sub-MDP $\mathcal{M}_{n}$. We employ a vector form for the rewards to offer the flexibility for formulating fairness objectives on individual expected total discounted rewards in later sections.
@@ 2.2 The Generalized Gini Function
The vector $\bm{V}^{\pi}_{0}$ represents the expected utilities for sub-MDPs. A social welfare function aggregates these utilities into a scalar, measuring fairness in utility distribution with respect to a maximization objective.
Social welfare functions can vary depending on the values of a society, such as $\alpha$-fairness (Mo and Walrand,, 2000), Nash social welfare (Fan et al.,, 2022; Mandal and Gan,, 2022), or max-min fairness (Bistritz et al.,, 2020; Cousins et al.,, 2022). Following Siddique et al., (2020), we require a fair solution to meet three properties: efficiency, impartiality, and equity. This motivates the use of GGF from economics (Weymark,, 1981), which satisfies these properties. For $N$ sub-MDPs, GGF is defined as $\operatorname{GGF}_{{\bm{w}}}[{\bm{v}}]:=\min_{\sigma\in{\mathbb{S}}^{N}}\sum_{n=1}^{N}{w}_{n}{v}_{\sigma(n)}$, where ${\bm{v}}\in{\mathbb{R}}^{N}$, ${\bm{w}}\in\Delta({\mathcal{N}})$ is non-increasing in $n$, i.e., $w_{1}\geq w_{2}\geq\cdots\geq w_{N}$. Intuitively, since $\operatorname{GGF}_{{\bm{w}}}[{\bm{v}}]=\sum_{n=1}^{N}{w}_{n}{v}_{\sigma^{*}(n)}$ with $\sigma^{*}$ as the minimizer, which reorders the terms of ${\bm{v}}$ from lowest to largest, it computes the weighted sum of ${\bm{v}}$ assigning larger weights to its lowest components. When the order of sub-MDPs is fixed, we use the equivalent formulation $\operatorname{GGF}_{{\bm{w}}}[{\bm{v}}]=\min_{\sigma\in\mathbb{S}^{N}}\sum_{n}w_{\sigma(n)}v_{n}$ as permuting either vector results in the same outcome.
As discussed in Siddique et al., (2020), GGF can reduce to special cases by setting its weights to specific values, including the maxmin egalitarian approach ($w_{1}\rightarrow 1,w_{2}\rightarrow 0,\ldots,w_{N}\rightarrow 0$) (Rawls,, 1971), regularized maxmin egalitarian ($w_{1}\rightarrow 1,w_{2}\rightarrow\epsilon,\ldots,w_{N}\rightarrow\epsilon$), leximin notion of fairness ($w_{k}/w_{k+1}\rightarrow\infty$) (Rawls,, 1971; Moulin,, 1991), and the utilitarian approach formally defined below for the later use in reducing the GGF problem.
@@
The utilitarian approach maximizes average utilities over all individuals but does not guarantee fairness in utility distribution, as some sub-MDPs may be disadvantaged to increase overall utility. The use of GGF offers flexibility by encoding various fairness criteria in a structured way. Moreover, $\operatorname{GGF}_{{\bm{w}}}[{\bm{v}}]$ is concave in ${\bm{v}}$, which has nice properties for problem reformulation.
@@ 2.3 The GGF-WCMDP Problem
We note that Lemma 3.1 in Siddique et al., (2020) establishes the optimality of stationary Markov policies for any multi-objective discounted infinite-horizon MDP under the GGF criterion. To obtain an optimal policy for the GGF-WCMDP problem (2), we introduce the following LP model with the GGF objective (GGF-LP):
@@
$\mathrm{s.t.}$ $\hskip-10.00002pt\lambda_{i}+\nu_{j}\leq w_{i}\sum\limits_{\bm{s}\in\mathcal{S}^{(N)}}\sum\limits_{\bm{a}\in\mathcal{A}^{(N)}}\bm{r}_{j}(\bm{s},\bm{a})q(\bm{s},\bm{a}),\,\forall i,j{\in{\mathcal{N}}},$ (3d)
$\hskip-20.00003pt\sum\limits_{\bm{a}\in\mathcal{A}^{(N)}}q(\bm{s},\bm{a})-\gamma\sum\limits_{\bm{s}^{\prime}\in\mathcal{S}^{(N)}}\sum\limits_{\bm{a}\in\mathcal{A}^{(N)}}q(\bm{s}^{\prime},\bm{a})p^{(N)}(\bm{s}|\bm{s}^{\prime},\bm{a})=\mu^{(N)}(\bm{s}),\quad\forall\bm{s}\in\mathcal{S}^{(N)},$
$\hskip-20.00003ptq(\bm{s},\bm{a})\geq 0,\quad\forall\bm{s}\in\mathcal{S}^{(N)},\quad\forall\bm{a}\in\mathcal{A}^{(N)}.$
@@
The dual form separates dynamics from rewards, with the expected discounted reward for sub-MDP $n$ given by $\sum_{\bm{s}\in\mathcal{S}^{(N)}}\sum_{\bm{a}\in\mathcal{A}^{(N)}}\bm{r}_{n}(\bm{s},\bm{a})q(\bm{s},\bm{a})$. The one-to-one mapping between the solution $q(\bm{s},\bm{a})$ and an optimal policy $\pi(\bm{s},\bm{a})$ is $\pi(\bm{s},\bm{a})=q(\bm{s},\bm{a})/\sum_{\bm{a}\in\mathcal{A}^{(N)}}q(\bm{s},\bm{a})$. Scalability is a critical challenge in obtaining exact solutions as the state and action spaces grow exponentially with respect to the number of sub-MDPs, making the problem intractable. We thus explore approaches that exploit symmetric problem structures, apply count-based state aggregation, and use RL-based approximation methods, to address this scalability issue, which will be discussed next.
@@ A WCMDP is symmetric if
This symmetry ensures that the expected state-value function, when averaged over all trajectories, is identical for each sub-MDP, leading to a uniform state-value representation. From this observation, and applying Theorem 6.9.1 from (Puterman,, 2005), we construct a permutation-invariant policy from any policy, resulting in uniform state-value representation (Lemma 3.3).
@@
This theorem simplifies solving the GGF-WCMDP problem by reducing it to an equivalent utilitarian problem, showing that at least one permutation-invariant policy is optimal for the original GGF-WCMDP problem and the utilitarian reduction. Therefore, we can restrict the search for optimal policies to this specific class of permutation-invariant policies. The utilitarian approach does not compromise the GGF optimality and allows us to leverage more efficient and scalable techniques to solve the GGF-WCMDP problem (Eq. 2), such as the Whittle index policies for RMABs, as demonstrated in the experimental section. We note that the utilitarian reduction theorem can be extended to a broader class of fairness measures, such as $\alpha$-fairness, as long as the fairness measure is concave, permutation invariant, and constant vector invariant (Corollary B.4.1). See Appendix B.5 for a detailed proof.